OpenStudy (anonymous):

42y^3 - 21y^2 ------------- 4y

OpenStudy (anonymous):

=(42/4)y^(3-1)-(21/4)y^(2-1) =(21/2)y^2 - (21/4)y =(21)y(1/2y - 1/4)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

you can simplify even more. Like this 21/4*y(2y-1)

OpenStudy (anonymous):

so the final answer is......not written as a fraction?

OpenStudy (anonymous):

no